OpenStudy (yrivers36):

For the function: f(x)= square root 4-2x a. plot the graph of the function and show points that were used. b. using interval notation give the domain and range of this function. c. give the x and y intercepts as ordered pairs. Please explain For the function: f(x)= square root 4-2x a. plot the graph of the function and show points that were used. b. using interval notation give the domain and range of this function. c. give the x and y intercepts as ordered pairs. Please explain @Mathematics

jhonyy9 (jhonyy9):

x can having values from all R without 2 - range equal 1

jhonyy9 (jhonyy9):

y=2 x=2

OpenStudy (eyust707):

\[f(x) = \sqrt{4-2x}\] a.  b. \[\left[ -\infty, 2 \right]\] c. y=2 x=2

OpenStudy (eyust707):

range includes o and goes to infinity